OpenStudy (anonymous):

HELP PLEASE!!! What is the 7th term of the geometric sequence where a1 = -3 and a5 = -768? 14. What is the 7th term of the geometric sequence where a1 = -4,096 and a4 = 64? 15. What is the sum of the geometric sequence -3, 21, -147, ... if there are 6 terms?

OpenStudy (campbell_st):

well a term in a geometric sqeuence ios \[a_{n} = a_{1} \times r^{n - 1}\] you need to find the common ratio so \[-768 = -3 \times r^{5 -1}...or..... -768 = -3 \times r^4\] solve for r then the sum is \[S_{n} = \frac{a_{1}(1 - r^n)}{1 - r}\] you ahve a1, you found r and know n = 7 just substitute to fine the sum similar for the other questions, find the common ratio r 1st